OpenStudy (anonymous):

Factor completely: 3ab(x + 1) − 2(x + 1)

OpenStudy (anonymous):

@ivettef365 @whpalmer4

OpenStudy (anonymous):

Please help!

OpenStudy (whpalmer4):

look at what you have - are there common factors between the two terms?

OpenStudy (anonymous):

3abx + 3ab?

OpenStudy (anonymous):

^not the answer

OpenStudy (whpalmer4):

no, how about (x+1)?

OpenStudy (anonymous):

wait what? Im confused already...

OpenStudy (whpalmer4):

it might be easier to see written this way: (x+1)3ab +(x+1)(-2)

OpenStudy (whpalmer4):

(I just reversed the order of the multiplications, which does not change the value)

OpenStudy (whpalmer4):

isn't that the same as (x+1)(3ab-2)?

OpenStudy (anonymous):

yesh

OpenStudy (anonymous):

3abx - 2x)(3ab - 2?

OpenStudy (anonymous):

I dont know what to do with the x....

OpenStudy (whpalmer4):

no, (x+1)(3ab-2) is the complete factoring or if you prefer (3ab-2)(x+1)

OpenStudy (anonymous):

wait wha? :-?

OpenStudy (whpalmer4):

you're just undoing the distributive property here... instead of going from \((p + q)(r + s) = p(r+s) + q(r+s) \) you are going the other way and taking \(p(r+s) + q(r+s) = (p+q)(r+s)\)

OpenStudy (anonymous):

oh ok

OpenStudy (whpalmer4):

it's sometimes a little hard to see the opportunities to do this at first, because you're not used to going in that direction...

OpenStudy (anonymous):

yeah...

OpenStudy (whpalmer4):

there's a technique called factoring by grouping that uses this...

OpenStudy (anonymous):

ok that makes sense now thx!

OpenStudy (whpalmer4):

great! you're welcome!